### v1

# Infrared and visible Image Fusion with Language-driven Loss in CLIP Embedding Space

Abstract
Infrared-visible image fusion (IVIF) has attracted much attention owing to the highly-complementary properties of the two image modalities. Due to the lack of ground-truth fused images, the fusion output of current deep-learning based methods heavily depends on the loss functions defined mathematically. As it is hard to well mathematically define the fused image without ground truth, the performance of existing fusion methods is limited. In this paper, we first propose to use natural language to express the objective of IVIF, which can avoid the explicit mathematical modeling of fusion output in current losses, and make full use of the advantage of language expression to improve the fusion performance. For this purpose, we present a comprehensive language-expressed fusion objective, and encode relevant texts into the multi-modal embedding space using CLIP. A language-driven fusion model is then constructed in the embedding space, by establishing the relationship among the embedded vectors to represent the fusion objective and input image modalities. Finally, a language-driven loss is derived to make the actual IVIF aligned with the embedded language-driven fusion model via supervised training. Experiments show that our method can obtain much better fusion results than existing techniques. The codes are available at ***.
11footnotetext: Corresponding author
1 Introduction
Multi-modal image fusion techniques, particularly the ones for infrared and visible image fusion (IVIF), have received considerable attention in recent yearsZhang and Demiris (2023). Infrared images reflect thermal radiation information, unaffected by lighting conditions, but often lack sufficient details. Conversely, visible images are often rich in details but sensitive to illumination, smoke and adverse weather. IVIF task aims to generate a fused image that integrate the complementary information from the two modalities. Currently, the deep-learning based methods dominate the field of IVIF. However, since there is no physical ground-truth fused image, the fusion output is greatly dependent on the loss functions defined mathematically for modelling the fused image.
Figure 1: We explore the potential of CLIP for infrared and visible image perception. The results show that CLIP has strong perception capability for both infrared and visible images. Furthermore, we investigate the impact of different language prompts on CLIP, and found that the prompts ‘a visible gray image’ and ‘an infrared image’ leads to higher classification scores for visible and infrared images, respectively.
Researchers have proposed to use various loss functions to supervise fusion network training, including contrastive lossZhu et al. (2022), adversarial lossRao et al. (2023); Ma et al. (2020), high-level task supervision lossXie et al. (2023); Tang et al. (2023), SSIM lossLi and Wu (2018); Li et al. (2023), etc. To improve the fusion results, current deep learning-based methods tend to combine various loss functions. Consequently, the overall loss function becomes increasingly complex. However, without knowing the real fused image, it is difficult for us to really well model the fused output mathematically with analytical loss functions. As a result, the performance of these fusion methods is limited.
Although it is hard to well model the fused image in mathematics, it is possible for us to seek for an alternative solution, i.e., using natural language to express the fusion objective. As we known, natural language can be used more easily to express the concept and intuitive ideas, while it would be much harder to convert them using precise mathematical symbols and formulations. Usually, before we conduct mathematical modeling for a problem, we need to be able to roughly describe the model through language, and then further transform it into mathematical formulas. Mathematical modeling, in contrast, is often much harder than language description, not to mention that many problems can not be explicitly modeled. Based on this principle, to solve the problem of IVIF, this paper first attempts to model the fusion objective based on natural language, and proposes a language-driven fusion loss to avoid complex and explicit mathematical modeling. We found that this approach is not only simpler, but also can achieve the fusion results much better than existing methods.
More specifically, to achieve the above goal, we try to leverage the currently popular pre-trained vision-language model CLIPRadford et al. (2021). As shown in Figure 1, we found that CLIP is able to perceive the infrared images, indicating that it can well encode the infrared image information (as that for the typical visible image) into the multi-modal embedding space of CLIP, in which the vision-language alignment has been achieved. To solve the issue of IVIF, we next establish an embedded language-driven fusion model using the text encoder of CLIP, for which we developed a language-expressed fusion objective, and use the prompts “an infrared image” and “a visible gray image” to describe the input infrared and visible images, respectively. Our goal is then become to make the fusion of infrared and visible images aligned with the language-driven fusion model in the CLIP embedding space through supervised training.
The main contributions are summarized as follows:
  • We first propose to use nature language to express the whole objective of IVIF, which allows to avoid the complex and explicit mathematical modeling in current fusion loss functions.
  • A language-driven fusion model is derived in the CLIP embedding space, based on which we develop a language-driven loss for IVIF and propose the first CLIP-based multi-modal image fusion method.
  • Experiments show a great improvement of fusion quality achieved by the proposed method, revealing the superiority of language in modeling of the fusion output and the potential of pre-trained vision-language model in improving the IVIF performance.
Figure 2: The Framework of the proposed method. The dashed line represents the language-driven training process, while the solid box shows the inference process.
2 Related Work
Infrared-visible Image Fusion Traditional IVIF methods utilize manually designed fusion rules and features to fuse infrared and visible information. These include multi-scale decomposition methods Vishwakarma and Bhuyan (2018); Zhou et al. (2016, 2023), sparse representation Wang et al. (2014), and saliency-based methods Han et al. (2013a); Ellmauthaler et al. (2012). In recent years, deep learning-based methods Ma et al. (2022); Cheng et al. (2023); Liu et al. (2021); Zhang et al. (2020) have become the mainstream approaches to address the inherent challenges of the IVIF task. Due to the lack of ground truth, the fusion result is highly dependent on the used loss functions. Approaches like U2Fusion Xu et al. (2020), DenseFuse Li and Wu (2018) and NestFuse Li et al. (2020) mainly utilize the pixel intensity loss and structure similarity Ma et al. (2015) loss to ensure that fusion results preserve the source image information. To obtain the fusion results with richer details, perceptual loss Long et al. (2021) are introduced for training. In addition, GAN-based methods Rao et al. (2023); Ma et al. (2020) use adversarial loss to preserve the salient thermal targets and texture details. More recent studies utilize high-level tasks, such as object detection Liu et al. (2022); Zhao et al. (2023), image classification Tang et al. (2022), and semantic segmentation Xie et al. (2023); Tang et al. (2023), to construct the semantic losses to improve the fusion performance. Although can help to obtain better object detection or segmentation performance, the fusion performance of these methods are constrained by a limited number of object or scene labels, and can not obtain a very comprehensive fused image.
Text-driven Image Generation with CLIP OpenAI proposed CLIP Radford et al. (2021), a high-performance text-image embedding model trained on 400 million text-image pairs. This model bridge the gap between image and language, establishing new benchmarks in related fields, including image generation. Techniques for generating and manipulating images driven by text increasingly utilize the CLIP model. Representative methods Patashnik et al. (2021); Gal et al. (2022) combines CLIP with StyleGAN Karras et al. (2019, 2020) to synthesize images by optimizing the latent code of StyleGAN, realizing image manipulation through text prompts. DiffusionCLIP Kim and Ye (2021) incorporates CLIP with a diffusion model Song and Ermon (2019) to enable text-baesd image manipulation, providing enhanced performance and more stable training compared with GAN-based approaches. So far, there is no literature to use CLIP to solve the challenge of multi-modal image fusion. For this purpose, a proper language-based fusion model should be established based on CLIP.
Figure 3: The overall schematics of the fusion process in CLIP embedding space. $\Delta V_{vs}$ and $\Delta V_{ir}$ jointly define the transition relationship established by the language-driven fusion model, while $\Delta\tilde{V}_{vs}$ and $\Delta\tilde{V}_{ir}$ denote the actual image fusion relationship.
3 Proposed Method
3.1 The Framework of Our Method
Figure 2 illustrates an overview of our method, which includes the language-driven training process and the inference process. In the language-driven training process, the infrared and visible images are input respectively to the image encoder of CLIP to obtain their embedded vectors in the CLIP embedding space. At the same time, the texts “an infrared image” and “a visible image” are encoded with the text encoder of CLIP to obtain their corresponding embedded vectors. Besides, we use another text to express the fusion result conceived in our mind and also produce its embedded vector with the text encoder. This enable us to construct an embedded language-driven fusion model in the CLIP embedding space, by establishing the relationship between the embedded vector from the text to express fusion objective and the embedded vectors from the texts to describe the different image modalities (will be detailed in Section 3.3).
Figure 4: The overall structure of the fusion network $G$.
Note that in the training process, the fused image obtained by the fusion network $G$ was also input to produce its embedded vector in the CLIP embedding space. The language-driven training of the fusion network $G$ is then become to align the relationship between the embedded vectors of input images and fused image with the fusion relationship established by the language-driven model. As the language-driven fusion model is not involved in the inference process, the actual fusion of images does not need to use CLIP, and thus no extra computation is introduced.
3.2 Language-expressed Fusion Objective
The CLIP model that can perceive both infrared and visible image information allows us to express the fusion objective using language, and then encode it into the CLIP embedding space. As illustrated in Figure 2, the language expression stems from our mind imaging what the fusion result should be. Using the language, it is possible for us to give a comprehensive description of the fusion effect with a short sentence, e.g., in this paper we describe it as “a vivid image with detailed background and obvious objects”. This expression can lead to produce the fused image with pretty rich and comprehensive information, covering both the background and various objects in the image. In contrast, it is difficult to use analytical loss functions to formulate this objective, because we can not define the “background” and “objects” with mathematical symbols, while the pre-trained language-vision model might have this cognition for an image. Some work Liu et al. (2022); Tang et al. (2023) leveraged the object detection or semantic segmentation results to constructed the supervision losses for IVIF. However, these approaches are constrained by a limited number of object or scene labels, whereas CLIP have the perception for various objects and backgrounds, not to mention that it can react to the words “obvious” and “detailed” in the objective description.
3.3 Language-driven Fusion Model and Loss
Embedded Language-driven Fusion Model. After having the language-expressed fusion objective encoded in the CLIP embedding space, the next issue is to establish a language-driven fusion model, which can guide to generate the described target image from the source images. As shown in Figure 3, let $V_{f}$ denote the embedded vector from the text to express the fusion objective, and $V_{vs}$, $V_{ir}$ denote the embedded vectors from the texts to describe the visible and infrared images, respectively. Inspired by the image generator with CLIP Gal et al. (2022), we construct the fusion model by establishing the transition relationship from the source vectors $V_{vs}$, $V_{ir}$ to the objective vector $V_{f}$. This results in the transition vectors $\Delta V_{vs}=V_{f}-V_{vs}$ and $\Delta V_{ir}=V_{f}-V_{ir}$ as illustrated by the dash arrows in Figure 3. The combination of these two vectors jointly defines the direction to generate the fused result from the input modalities in the embedding space.
Language-driven Fusion Loss. To obtain the fused image as the language described, we make the actual fusion of infrared and visible images aligned with the embedded language-driven fusion model during the training. For this purpose, the source images and the fused image obtained by the fusion network are input to the image encoder of CLIP (see Figure 2), and obtain the image encoded vectors $\tilde{V}_{vs}$, $\tilde{V}_{ir}$ and $\tilde{V}_{f}$, respectively. As illustrated in Figure 3, due to the alignment of vision-language achieved by CLIP, the image-text encoded vector pairs $\tilde{V}_{vs}\sim V_{vs}$, $\tilde{V}_{ir}\sim V_{ir}$ can be considered to share the same locations in the embedding space, respectively. Thus, to make the actual image fusion aligned with the language-driven model, we need only to force the image-encoded transition vectors $\Delta\tilde{V}_{vs}$ ($\Delta\tilde{V}_{vs}=\tilde{V}_{f}-\tilde{V}_{vs}$) and $\Delta\tilde{V}_{ir}$ ($\Delta\tilde{V}_{ir}=\tilde{V}_{f}-\tilde{V}_{ir}$) parallel to the text-encoded transition vectors $\Delta\tilde{V}_{vs}$ and $\Delta\tilde{V}_{ir}$, respectively. In this case, as illustrated by the two solid arrows in Figure 3, it would drive the fusion towards the language-described direction and reach to the objective state in the embedding space. As a result, we define a fusion direction loss as follows:
[…]
The above loss ensures that $\Delta\tilde{V}_{vs}$ and $\Delta\tilde{V}_{ir}$ are parallel to $\Delta{V}_{vs}$ and $\Delta{V}_{ir}$, respectively. $\mathcal{L}_{d}$ is performed on the entire image. In order to ensure the local fusion quality, in practice we also use the losses performed on multi-scale patches. In our implementation, we achieve this by randomly selecting a number of patches with the sizes of 128×128 and 64×64 in the images. This leads to an enhanced fusion direction loss $\mathcal{L}_{d}^{\dagger}$ taking into account both local and global performance. In addition, to improve the robustness and generalization of the fusion model, we also introduce a regularization term into the loss, and ultimately obtain a language-driven fusion loss as:
$\begin{gathered}\mathcal{L}_{lg}=\mathcal{L}_{d}^{\dagger}+\lambda\Phi,\\ \end{gathered}$ (6)
where $\Phi$ is the regularization term, and $\lambda$ denotes its weight ($\lambda=0.5$ is typically used). As shown in Figure 3, the training of fusion model in the embedding space should prevent the embedded vector of fused image $\tilde{V}_{f}$ being identical to any of the source vectors $\tilde{V}_{vs}$ and $\tilde{V}_{ir}$. This can be achieved by simultaneously imposing penalty on the angles $\theta_{vs}$ (between $\tilde{V}_{f}$ and $\tilde{V}_{vs}$) and $\theta_{ir}$ (between $\tilde{V}_{f}$ and $\tilde{V}_{ir}$). To this end, the regularization term $\Phi$ can be defined as:
$\begin{gathered}\Phi=|1-\frac{\tilde{V}_{f}\cdot\tilde{V}_{vs}}{\|\tilde{V}_{f}\|\cdot\|\tilde{V}_{vs}\|}|+|1-\frac{\tilde{V}_{f}\cdot\tilde{V}_{ir}}{\|\tilde{V}_{f}\|\cdot\|\tilde{V}_{ir}\|}|.\\ \end{gathered}$ (9)
The tendency of $\tilde{V}_{f}$ becoming parallel to any of the source vectors would make the other angle ($\theta_{vs}$ or $\theta_{ir}$) become large, which can be suppressed by using $\Phi$.
Finally, we need to introduce a feature-fidelity loss $\mathcal{L}_{v}$ to suppress unnecessary “imaged” content generated by the language-driven model. It is defined as follows:
$\begin{gathered}\mathcal{L}_{v}=\sum(\|\psi_{f}^{k}-max(\psi_{ir}^{k},\psi_{vs}^{k})\|),\\ \end{gathered}$ (12)
where $\psi_{f}^{k}$, $\psi_{ir}^{k}$ and $\psi_{vs}^{k}$ are the feature maps extracted by the $k$-th convolution module in the pre-trained VGG-19 network, with the input of the fused, infrared and visible images, respectively. $k$ is typically in the set ${5,10,21}$. Therefore, our total loss function consists of the language-driven fusion loss and feature-fidelity loss, formulated as:
$\begin{gathered}\mathcal{L}_{total}=\mathcal{L}_{lg}+\alpha\mathcal{L}_{v}.\\ \end{gathered}$ (15)
Figure 5: The detailed structure of the cross fusion attention.
Figure 6: Comparison of the fusion results from different methods on the images in TNO dataset.
FusionGAN MFEIF PIAFusion PMGI RFN SwinFusion U2Fusion UMF GANMcC Ours
EN 6.550 6.749 6.929 7.058 7.086 6.908 7.035 6.629 6.791 7.335
AG 3.069 3.685 6.029 4.616 3.066 5.801 6.430 4.113 3.395 9.878
SD 30.487 33.827 41.400 38.707 40.224 39.735 37.894 30.817 34.162 51.502
SF 3.922 4.345 6.291 5.232 3.837 6.166 6.787 4.674 4.082 8.365
VIFF 0.265 0.376 0.405 0.593 0.575 0.451 0.699 0.359 0.433 0.751
Table 1: Quantitative assessment using five metrics (for TNO dataset), with the best results highlighted in red, the second best highlighted in blue.
3.4 Fusion Network
The architecture of our infrared-visible fusion network is illustrated in Figure 4. It consists of three main components. Initially, the two-branch encoder is designed to independently extract features from visible and infrared images, ensuring that information of two modalities are captured. We then introduce a cross fusion attention used to fuse extracted features from the two-branch encoder. In the final stage of image reconstruction, we utilize a simple decoder to generate fusion results, which consist of a feature-enhancement module and four convolution layers with layer normalization and ReLU activation.
[…]
$S_{vs}$ $=\text{softmax}\left(\frac{Q_{vs}K_{vs}^{T}+W_{vs}}{\sqrt{d_{k}}}\right),$ (16)
$S_{ir}$ $=\text{softmax}\left(\frac{Q_{ir}K_{ir}^{T}+W_{ir}}{\sqrt{d_{k}}}\right),$ (17)
$F_{a}$ $=\text{reshape}\left(S_{vs}V\|S_{ir}V\right).$ (18)
Figure 7: Comparison of the fusion results from different methods on the images in RoadScene dataset.
Figure 8: Comparison of the ablation study on language-driven fusion loss and cross fusion attention.
4 Experiments
FusionGAN MFEIF PIAFusion PMGI RFN SwinFusion U2Fusion UMF GANMcC Ours
EN 7.073 7.042 6.908 7.328 7.337 6.877 7.263 7.047 7.237 7.474
AG 3.911 4.329 5.191 5.296 3.529 5.250 7.298 4.888 4.312 10.422
SD 39.151 39.264 42.957 46.211 46.055 44.255 42.918 37.277 43.836 53.481
SF 5.019 5.288 5.837 6.006 4.708 5.956 7.207 5.651 5.317 8.059
VIFF 0.305 0.414 0.353 0.467 0.479 0.442 0.550 0.410 0.454 0.556
Table 2: Quantitative assessment using five metrics (for RoadScene dataset), with the best results highlighted in red, the second best highlighted in blue.
4.1 setup
Datasets. We conduct experiments on three public datasets: M3FD Liu et al. (2022), RoadScene Xu et al. (2020), and TNO Toet (2017). We use M3FD for network training, as it covers various scenes, illumination, and weather conditions. The TNO dataset includes multi-spectral, military-relevant images, while RoadScene focuses on outdoor scenes. The TNO and RoadScene datasets are used to evaluate the performance of our proposed method.
Compared Methods. We compare our method with nine state-of-the-art fusion methods, including FusionGANMa et al. (2019), MFEIFLiu et al. (2021), PIAFusionTang et al. (2022), PMGIZhang et al. (2020), RFNLi et al. (2021), SwinFusionMa et al. (2022), U2FusionXu et al. (2020), UMFDi et al. (2022) and GANMcCMa et al. (2021).
Metrics. Five metrics are used to evaluate our method: information entropy (EN)Roberts et al. (2008), average gradient (AG)Cui et al. (2015), standard deviation (SD)Ma et al. (2019), spatial frequency (SF)Roberts et al. (2008) and the visual information fidelity for fusion (VIFF)Han et al. (2013b). EN measures the amount of information contained in the fusion result. AG measures image gradients, with higher values signifying richer detail. SD is used to quantify the contrast in images. SF evaluates the level of detail in an image. VIFF evaluates the quality of image fusion, considering how human vision perceives these images.
Implementation Details. Our framework is implemented with PyTorch on two NVIDIA RTX 3080 GPUs. we employ the Adam optimizer with a learning rate set to 0.0003. The training process consists of 100 epochs, with a warm-up phase during the first epoch. After 50 epochs, a decay strategy is implemented, halving the learning rate every 10 epochs. $\alpha$ is set to 1 for training. The raw data is randomly cropped into 256x256 pixel images with batches of 12, which are used as inputs for the fusion network.
4.2 Comparison with State-of-the-art Methods
qualitative comparison We compare our method with nine other state-of-the-art (SOTA) methods to demonstrate the superiority of our method, as shown in Figure 6 and Figure 7. As can be seen, our method generates much better fusion results compared to others. The fused images generated by our method have four significant advantages: firstly, the fusion results have superior visual perception and natural contrast; secondly, the objects in fused images (e.g., cars and buildings marked by red boxes) are more salient; thirdly, the texture details in the background are better preserved; lastly, our method can overcome the effects of low-quality source images and generates fusion results that are consistent with the language-expressed fusion objective.
In contrast, the fused images of MFEIF, PIAFusion, SwinFUsion, and UMF have suboptimal contrast and visual effect, and the objects in their fusion results are not salient (as shown by the red boxes in Figure 6). FusionGAN, PMGI, RFN, U2Fusion and GANMcC obtain contrasted images, but they do not effectively fuse important information in the visible image, resulting in darker sky and blurrier objects (see the red box in Figure 6). Our method obtains much superior fusion results with overally better visual perception, clearer targets and richer background information. In addition, we also evaluate the performance of our method in low-light environment at night, as shown in Figure 7. Our approach better integrates the details from source images and highlights the objects in fused images, as demonstrated by the red car in Figure 7. Besides, our method is able to significantly enhance background details, such as building structures, trees, etc, as can be seen in the red box of Figure 7.
quantitative comparison Tables 1 and 2 summarize the performance comparison of our method with nine state-of-the-art (SOTA) methods on five metrics. On the metrics EN, AG, SD and SF, our method obtains the best performance both on the TNO dataset and the RoadScene dataset. Among these metrics, the highest EN value indicates that our method is able to fuse more information from the source images, while the performance of SD and VIFF proves that our method have natural contrast and better perceptual quality . The highest scores of AG and SF further prove the richness of our method in terms of detail information.
4.3 Ablation Study
Language-driven Fusion Loss. Figure 8 illustrates the results of our ablation study. Figure 8(c) and Figure 8(e) show the fused images obtained from the method with and without the language-driven fusion loss (LDL), respectively. It can be seen that the results without LDL have poor performance on overall visual perception, contrast and clarity of details, while the method with LDL generates much better results, even if the quality of source images is not good.
From the fusion results in the top row, it can be seen that the pedestrian is not obvious and clear in red box of Figure 8(c), which indicates the method without LDL cannot fully fuse salient information from source images. In contrast, objects in the fused images of the method with LDL always have higher salience and clarity as can be seen from the people in the red box. In the second row, the fusion result of the method without LDL is unsatisfactory in terms of both background and objects, especially the car in the red box. Moreover, the method with LDL can also enhance the details as can be seen from the red box in the last row. The window structure of Figure 8(e) is much clearer than Figure8(c), and thermal radiation information from infrared image is also preserved better.
Cross Fusion Attention. Figure 8(d) and Figure 8(e) illustrate the results with and without the cross fusion attention (CFA), respectively. In the first row, the image of Figure 8(d) exhibits rough edges (see the person in the red box), while the results of our method have better performance on the details and structures. It can be seen that the structure of car (as shown in the red box) in Figure 8(e) is more salient than Figure 8(d). The method without CFA have suboptimal fusion performance on background, and from the second row, we can find that it can not well fuse the information from sky. The lack of CFA results in local darkness (see the sky in third row). From the window in the red boxes, it can be seen that the method without CFA can not effectively fuse the multi-modal information for the window in the image. In contrast, our method obtain better fusion performance on background and objects in fused image.
5 Conclusion
In this paper, we use natural language to express the whole objective of infrared-visible image fusion, thereby avoiding complex and explicit mathematical modeling in the fusion loss functions. To achieve this goal, we construct an embedded language-driven fusion model based on the CLIP text encoder. This model defines the expected fusion directions within the CLIP embedding space, which are used to guide the fusion network better to fuse source images. To generate the fusion results consistent with the fusion objective expressed through language, we next derive a language-driven fusion loss based on the model. This loss is employed to align the actual fusion of infrared and visible images to the embedded language-driven fusion model through supervised training process. By minimizing this loss, the fusion network can successfully generate the fused image as the language described. Experimental results demonstrate that our method significantly improves the quality of fused images.